OpenStudy (anonymous):

A set of equations is given below: Equation A: g = 4h + 1 Equation B: g = 3h + 5 Which of the following is a step that can be used to find the solution to the set of equations? 4h + 5 = 3h + 1 4h + 1 = 3h + 5 4h = 3h + 5 4h + 1 = 3h

OpenStudy (anonymous):

can you help me please

OpenStudy (campbell_st):

both equations are equal to g to that means that you can write them with the right side of equation A equal to the right side of equation B hope it helps

OpenStudy (anonymous):

can you help with this one too Rachel is 11 years older than Scott. Rachel's age is 7 years less than two times Scott's age. The system below models the relationship between Rachel's age (r) and Scott's age (s): r = s + 11 r = 2s – 7 Which is the correct method to find Rachel's and Scott's ages? Solve r + 11 = 2r – 7 to find the value of s. Solve s + 11 = 2s – 7 to find the value of s. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (campbell_st):

same thing, they are both equal to r so right side of equation 1 is equal to right side of equation 2

OpenStudy (campbell_st):

you may also tell you teacher that you can't find where the graphs intersect the x-axis or y-axis as the equations are written in terms of r and s

OpenStudy (anonymous):

i don't get what the answer will be